OpenStudy (ttop0816):

please help! What is the equation of the line that passes through (-2, 2) and (1, -4)? -y = 2x - 2 -y = -2x - 4 -y = -2x - 2 -y = 2x + 4

OpenStudy (uri):

First find slope Slope formula:\[\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (anonymous):

thats the formula hahahaha correct :D

OpenStudy (uri):

(-2,2) (1,-4) x1,y1 x2,y2 Plug in the values and solve @ttop0816

OpenStudy (ttop0816):

@jacalneaila @uri thank you so much!

OpenStudy (anonymous):

ok sorry uri

terenzreignz (terenzreignz):

lol @uri go easy on her And yes @ttop0816 that's just the slope... which is only the beginning :D

OpenStudy (anonymous):

this is the guide @ttop0816

terenzreignz (terenzreignz):

The website doesn't load @jacalneaila ^_^

OpenStudy (ttop0816):

oh.... haha please explain the rest please!

terenzreignz (terenzreignz):

Gladly @ttop0816 But first, please find the slope of the line. I think it was briefly mentioned...

OpenStudy (anonymous):

ah really

OpenStudy (anonymous):

?why? lol XD

terenzreignz (terenzreignz):

Hey @uri please be nice :(

OpenStudy (ttop0816):

@uri okay i found the slope

OpenStudy (ttop0816):

@uri -2

OpenStudy (anonymous):

@uri be calm

terenzreignz (terenzreignz):

Brilliant :) When you know the slope \(\large \color{blue}m\) of a line as well as a point \(\large (\color{red}a \ , \ \color{green}b)\) which the line passes through, then the equation of the line is given by \[\Large y = \color{blue}m (x-\color{red}a)+\color{green}b\]

terenzreignz (terenzreignz):

Savvy? XD

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@terenzreignz :D

terenzreignz (terenzreignz):

^_^ may pagkataray si uri

OpenStudy (ttop0816):

can anyone tell me the answer or how to solve the question????

terenzreignz (terenzreignz):

I am... you know the slope m = -2 right? Now, what's a point that your line passes through... you're given two of them, we only need one, pick one ^_^

OpenStudy (ttop0816):

@terenzreignz would it be this one?? --->-y = 2x - 2

terenzreignz (terenzreignz):

I don't know @ttop0816 let's find out. Now, pick one of your points...

OpenStudy (uri):

terence what formula is this....havent used it ever.

OpenStudy (uri):

is this same as y-y1=m(x-x1)?

OpenStudy (ttop0816):

@terenzreignz @uri im sorry but could i just get the answer this one time please? im in a hurry right now;

terenzreignz (terenzreignz):

@uri's decision...

OpenStudy (uri):

First tell me..Is the formula right? @terenzreignz

terenzreignz (terenzreignz):

Of course @uri It comes from the point slope form \[\Large y-b = m(x-a)\]or if you like \[\Large y - y_0 = m(x-x_0)\] I just brought the b-over to the right side.

OpenStudy (uri):

Why did you write y0 it's y1

terenzreignz (terenzreignz):

Doesn't really matter :)

OpenStudy (uri):

but aren't you confusing this kid? Because at schoool we usually use y-y1=m(x-x1) @terenzreignz

terenzreignz (terenzreignz):

Yes, we use that, but if you notice, it looks NOTHING like her choices.

terenzreignz (terenzreignz):

I adapt to the one that will most easily be recognisable.

OpenStudy (uri):

Oh..so help her out in this one..@terenzreignz

terenzreignz (terenzreignz):

I'm willing. But... work with me here @ttop0816 Come on, I'm just asking you (for now) to pick one of your two points...